# RoboCOIN: An Open-Sourced Bimanual Robotic Data COllection for INtegrated Manipulation

### v1

@@ II Related Work
Figure 2: Data collection platforms of our RoboCOIN. (a) Dual-Arm Robot (e.g., Agilex Cobot Magic). A bimanual robot with two 6-DoF arms and parallel grippers, capable of performing complex bimanual tasks. (b) Half-Humanoid Robot (e.g., Realman RMC-AIDA-L). Left: wearable teleoperation device for human demonstration, right: bimanual robot with parallel grippers. (c) Humanoid Robot (e.g., Unitree G1edu-u3). Left: wearable motion capture device for human demonstration, right: humanoid robot with advanced dexterous manipulation capabilities. Head and wrist cameras provide multi-view visual observations.
@@ III RoboCOIN Dataset
Figure 4: The RoboCOIN framework introduces a hierarchical capability pyramid, structured across three levels: (a) trajectory-level annotations define the global concepts and task objectives; (b) segment-level annotations decompose the task into executable subtasks; and (c) frame-level annotations provide dense low-level details such as motion trajectories and gripper states. All annotations are temporally synchronized to form a cohesive data structure.
@@ III-A Data Collection and Storage
The RoboCOIN framework leverages a diverse set of 15 robotic platforms for comprehensive data acquisition, encompassing bimanual, half-humanoid, and humanoid configurations. Figure 2 illustrates three representative platforms: bimanual robots (e.g., Agilex Cobot Magic), half-humanoid robots (e.g., Realman RMC-AIDA-L), and humanoid robots (e.g., Unitree G1edu-u3). The framework employs teleoperation to ensure high-quality data collection, utilizing methods such as (a) leader-follower isomorphic arms, (b) exoskeletons, and (c) motion capture systems. The complete list of robotic platforms is detailed in Table 1. The platforms are equipped with a comprehensive suite of sensors. These capture multimodal data streams (RGB and depth) from multiple camera views (e.g., head, wrist, third-person, chest, and back), along with the robot’s kinematic state (including joint angles, end-effector poses, and gripper articulation). Essential environmental parameters, such as platform elevation and workspace geometry, are also recorded. All kinematic measurements adhere to standardized conventions: distances are in meters, end-effector orientations are represented by 6D rotation matrices within a unified left-handed coordinate system, and the gripper state is normalized to a continuous value from 0 to 1. Strict temporal synchronization across all data streams is maintained through timestamp alignment, ensuring consistency between visual observations and kinematic states.
@@ III-C Hierarchical Capability Pyramid
Trajectory-level Concepts. Trajectory-level annotations represent the concepts of a task by describing the overall scene configuration. This includes scene description (environment settings, object placements), and detailed attributes (e.g., color, shape, material, texture, and size). The resulting integrated representation supports global spatial and physical reasoning across the task sequence.
@@ V Experiments and Analysis
In this section, we evaluate the effectiveness of the RoboCOIN dataset and the CoRobot framework. This study aims to answer the following key research questions (RQs) through a series of experiments in real-world environments:
@@ V-A Experiment Setup
  • Unitree G1edu-u3[unitreeg1], a humanoid robot with two 7-DoF arms and dexterous hands, with full-body mobility.
@@ V-B Impact of Hierarchical Annotation Integration
Figure 8: Impact of RTML on data quality and model performance. (a) RTML Filtering. The amount of data removed by RTML in the two tasks. (b) Phase-wise and Metric-wise Analyses. The former identifies failure-prone operational stages, while the latter pinpoints frequently violated constraints. (c) Unitree G1edu-u3 + GR00T-N1.5. The performance of the two tasks across the four fine-tuning settings: GR00T-Raw, GR00T-Coarse, GR00T-Fine, and GR00T-Mine, as measured by success rate.
@@ V-C Impact of RTML
To address RQ3, we conducted experiments on two tasks using the Unitree G1edu-u3 platform: a single-arm task "pick the grape and place into the plate" and a bimanual task "push the bowl and place bread pieces into it". As shown in Figure 8(a), RTML filtering removed an average of 35.3% of low-quality trajectories across both tasks, indicating that a substantial portion of human demonstrations contains inconsistencies that could adversely affect policy learning. As shown in Figure 8(a) (b), we analyzed trajectories from two tasks using two complementary approaches: phase-wise and metric-wise analyses. The former identified failure-prone operational stages (e.g., move, grasp, place, etc.) by assessing segments against phase-specific constraints, while the latter pinpointed frequently violated constraints (e.g., workspace, velocity, acceleration, etc.) through evaluation against individual metrics. The phase-wise analysis revealed that the majority of trajectory failures occurred during the grasping phase, accounting for 52.7% of disqualifications, followed by the moving phase at 17.8%. On the other hand, the metric-wise analysis indicated that velocity violations were the most common cause of trajectory disqualification, responsible for 46.2% of failures, followed by duration violations at 24.5%. The complementary insights from both analyses highlight critical areas for improving demonstration quality, particularly in enhancing grasping techniques and adhering to velocity constraints during task execution.

### v2

@@ II Related Work
Figure 2: Data collection platforms of our RoboCOIN. (a) Dual Robot (e.g., Agilex Cobot Magic). A bimanual robot with two 6-DoF arms and parallel grippers, capable of performing complex bimanual tasks. (b) Half-Humanoid Robot (e.g., Realman RMC-AIDA-L). Left: wearable teleoperation device for human demonstration, right: bimanual robot with parallel grippers. (c) Humanoid Robot (e.g., Unitree G1edu). Left: wearable motion capture device for human demonstration, right: humanoid robot with advanced dexterous manipulation capabilities. Head and wrist cameras provide multi-view visual observations.
@@ III RoboCOIN Dataset
Figure 4: The RoboCOIN framework introduces a hierarchical capability pyramid, structured across three levels: (a) trajectory-level annotations define the global concepts; (b) segment-level annotations decompose the task into executable subtasks; and (c) frame-level annotations provide dense low-level details such as motion trajectories and gripper states. All annotations are temporally synchronized to form a cohesive data structure.
@@ III-A Data Collection and Storage
The RoboCOIN framework leverages a diverse set of 15 robotic platforms for comprehensive data acquisition, encompassing bimanual, half-humanoid, and humanoid configurations. Figure 2 illustrates three representative platforms: bimanual robots (e.g., Agilex Cobot Magic), half-humanoid robots (e.g., Realman RMC-AIDA-L), and humanoid robots (e.g., Unitree G1edu). The framework employs teleoperation to ensure high-quality data collection, utilizing methods such as (a) leader-follower isomorphic arms, (b) exoskeletons, and (c) motion capture systems. The complete list of robotic platforms is detailed in Table 1. The platforms are equipped with a comprehensive suite of sensors. These capture multimodal data streams (RGB and depth) from multiple camera views (e.g., head, wrist, third-person, chest, and back), along with the robot’s kinematic state (including joint angles, end-effector poses, and gripper articulation). Essential environmental parameters, such as platform elevation and workspace geometry, are also recorded. All kinematic measurements adhere to standardized conventions to ensure consistency across diverse robotic platforms. The joint state is measured in radians. The end-effector state comprises its position and orientation: distances are expressed in meters, and the orientation is represented using Euler angles (in radians) within a unified right-handed coordinate system (x-forward, y-left, z-up). The gripper state is normalized to a continuous value from 0 to 1. Strict temporal synchronization across all data streams is maintained through timestamp alignment, ensuring consistency between visual observations and kinematic states.
@@ III-C Hierarchical Capability Pyramid
Trajectory-level Concepts. Trajectory-level annotations capture task concepts by characterizing the overall scene configuration. This involves a description of the scene, detailing the environment settings and the placement of objects. The resulting unified representation enables global spatial and physical reasoning throughout the entire task sequence.
@@ V Experiments and Analysis
In this section, we evaluate the effectiveness of the RoboCOIN dataset and the CoRobot framework, focusing on multi-embodiment adaptability, the impact of hierarchical capability pyramid, and the RTML-validated data quality. This study aims to answer the following key research questions (RQs) through a series of experiments in real-world environments:
@@ V-A Experiment Setup
  • Unitree G1[unitreeg1], a humanoid robot with two 7-DoF arms and dexterous hands, with full-body mobility.
@@ V-B Impact of Hierarchical Annotation Integration
Figure 8: Impact of RTML on data quality and model performance. (a) RTML Filtering. The amount of data removed by RTML in the two tasks. (b) Phase-wise and Metric-wise Analyses. The former identifies failure-prone operational stages, while the latter pinpoints frequently violated constraints. (c) Unitree G1 + GR00T-N1.5. The performance of the two tasks across the four fine-tuning settings: GR00T-Raw, GR00T-Coarse, GR00T-Fine, and GR00T-Mine, as measured by success rate.
@@ V-C Impact of RTML
To address RQ3, we conducted experiments on two tasks using the Unitree G1 platform: a single-arm task "pick the grape and place into the plate" and a bimanual task "push the bowl and place bread pieces into it". As shown in Figure 8(a), RTML filtering removed an average of 35.3% of low-quality trajectories across both tasks, indicating that a substantial portion of human demonstrations contains inconsistencies that could adversely affect policy learning. As shown in Figure 8(a) (b), we analyzed trajectories from two tasks using two complementary approaches: phase-wise and metric-wise analyses. The former identified failure-prone operational stages (e.g., move, grasp, place, etc.) by assessing segments against phase-specific constraints, while the latter pinpointed frequently violated constraints (e.g., workspace, velocity, acceleration, etc.) through evaluation against individual metrics. The phase-wise analysis revealed that the majority of trajectory failures occurred during the grasping phase, accounting for 52.7% of disqualifications, followed by the moving phase at 17.8%. On the other hand, the metric-wise analysis indicated that velocity violations were the most common cause of trajectory disqualification, responsible for 46.2% of failures, followed by duration violations at 24.5%. The complementary insights from both analyses highlight critical areas for improving demonstration quality, particularly in enhancing grasping techniques and adhering to velocity constraints during task execution.